# Finding Maximum Independent Sets in Dynamic Graphs using Unsupervised Learning

### v1

@@ Abstract
We present the first unsupervised learning model for finding Maximum Independent Sets (MaxIS) in dynamic graphs where edges change over time. Our method combines structural learning from graph neural networks (GNNs) with a learned distributed update mechanism that, given an edge addition or deletion event, modifies nodes’ internal memories and infers their MaxIS membership in a single, parallel step. We parameterize our model by the update mechanism’s radius and investigate the resulting performance–runtime tradeoffs for various dynamic graph topologies. We evaluate our model against state-of-the-art MaxIS methods for static graphs, including a mixed integer programming solver, deterministic rule-based algorithms, and a heuristic learning framework based on dynamic programming and GNNs. Across synthetic and real-world dynamic graphs of 100–10,000 nodes, our model achieves competitive approximation ratios with excellent scalability; on large graphs, it significantly outperforms the state-of-the-art heuristic learning framework in solution quality, runtime, and memory usage. Our model generalizes well on graphs 100x larger than the ones used for training, achieving performance at par with both a greedy technique and a commercial mixed integer programming solver while running 1.5–23x faster than greedy.
@@ 1 Introduction
Combinatorial optimization problems on graphs (e.g., Graph-Coloring, Traveling-Salesman, Vehicle-Routing, etc.) arise naturally in many practical domains (Applegate et al., 2007; Toth and Vigo, 2014; Hochbaum, 1997). However, many of these problems are classically intractable to solve exactly, and some even resist efficient approximation. The situation is exacerbated by the fact that many real-world applications are dynamic, requiring algorithms to not only find solutions but also maintain them as the underlying structure evolves. While one could iteratively apply algorithms designed for static graphs at each time step, this approach is often computationally expensive and inefficient. Critically, this approach ignores the often close relationship between successive graph structures. A more effective strategy might avoid repeatedly recomputing solutions from scratch by instead leveraging these small structural changes to incrementally update an existing solution. Such update algorithms are an active area of theoretical algorithms research (Hanauer et al., 2022; Chechik and Zhang, 2021; Behnezhad et al., 2019), but no learning algorithms have yet taken this approach.
@@
In this paper, we consider the Maximum-Independent-Set (MaxIS) problem in a dynamic setting. For static graphs, MaxIS is classically known to be NP-hard (Garey and Johnson, 1979) and hard to approximate: for general graphs, deterministic approximation within a constant factor is impossible (Robson, 1986) and within an $n^{1-\varepsilon}$ factor is NP-hard (Håstad, 1996). Nevertheless, MaxIS has a diverse range of real-world applications, from identifying functional nodes in brain networks (Afek et al., 2011) to constructing diversified investment portfolios (Hidaka et al., 2023) to scheduling in public transportation systems (Damir N. et al., 2018). Each use case can be modeled as a conflict graph, where an edge represents a conflict between selecting two nodes. For example, in investment portfolio construction, diversification is critical; if one stock declines, others in the portfolio should not be closely related. In this case, the nodes represent investments, and the edges represent dependencies or correlations between them. Similarly, in transportation scheduling, conflicts may arise due to overlapping routes or timing constraints. In all these situations, the conflict graph changes dynamically due to factors such as stock market fluctuations, train delays, and other unpredictable events. Our goal is to solve MaxIS on such dynamically evolving graphs.
@@
Recently, the learning community has developed several methods for solving combinatorial optimization problems on static graphs (Bengio et al., 2021). Central to many of these methods are Graph Neural Networks (GNNs) due to their efficiency in capturing the structural information present in the real-world networks such as the Internet, social networks, and molecular interactions (Wu et al., 2021; Zhang et al., 2019). In the dynamic setting, recent models such as temporal GNNs are primarily concerned with tasks such as edge prediction, node classification, graph classification, etc. (Rossi et al., 2020; Xu et al., 2019). Work considering combinatorial optimization problems in a dynamic setting remains limited, with few examples such as Gunarathna et al. (2022), which is an indirect heuristic learning approach. However, no models exist to learn update mechanisms comparable to theoretical, rule-based update algorithms.
@@
We propose a model that solves MaxIS in dynamic graphs by directly learning an update mechanism instead of a heuristic learning. We utilize the power of message-passing GNNs (Battaglia et al., 2018) combined with sequence learning to learn an update mechanism that maintains an approximate MaxIS solution as the underlying graph changes. GNNs learn to aggregate structural information and update a node’s membership in the MaxIS solution at any instance of a dynamic graph, while sequence learning learns the effect of a change in the dynamic graph on the change in the MaxIS solution. Having learned an update mechanism, the model is able to update a MaxIS solution following a change in the dynamic graph in a single inference step, unlike more popular heuristic learning methods which require multiple inference steps and thus scale poorly with graph size. We underscore this as a necessary shift in perspective required to address some of the common criticisms of learning approaches to solving combinatorial optimization problems, namely, lack of scalability, computation inefficiency, and huge training sample requirements.
@@
We present the first unsupervised learning model for solving the Maximum-Independent-Set (MaxIS) problem on dynamic graphs. Compared to a commercial mixed integer programming solver (Gurobi Optimization, LLC, 2024), our model maintains competitive approximation ratios across a variety of dynamic graph topologies, sizes, and edge dynamics. For larger datasets, the model conclusively outperforms DP-GNN (Brusca et al., 2024), the state-of-the-art GNN-based model for MaxIS in static graphs. Resource scaling for our method vs. DP-GNN is drastic; we obtain solutions up to $10^{4}$ times faster and use up to 27.5x less memory. Our model is also the first learning method to successfully learn an update mechanism for the dynamic version of any NP-hard combinatorial optimization problem.
@@ 2 Related Work
Maximum Independent Set (MaxIS), an NP-complete problem, has been a long-standing problem in graph theory and has been explored since the 1970s in the field of combinatorial optimization. It has been studied extensively in the context of static graphs. There have been a number of attempts to achieve an exact algorithm for the problem using reduction rules (Tarjan and Trojanowski, 1977), branching heuristics (Fomin and Kaski, 2013), and kernelization techniques (Lamm et al., 2017) that compress the graph size being worked on. Xiao and Nagamochi (2017) have an algorithm that promises a run time of $1.1996^{n}n^{O(1)}$ and polynomial space where $n$ is the number of vertices in the graph.
While a wide variety of work exists with respect to static graphs, the dynamic graph instances relatively remain unexplored. Using the rule-based optimization techniques, the dynamic network version of the problem has been explored in the past 5 years. Zheng et al. (2019) and Gao et al. (2022) explore how these techniques can be exploited to maintain a MaxIS for a dynamically evolving network. These approaches typically rely on maintaining and updating an existing solution incrementally in response to edge insertions or deletions.
@@
We primarily look at the dynamic graph models called memory-based models due to their efficiency and success in learning prediction and classification tasks in dynamic graphs (Rossi et al., 2020; Kumar et al., 2019; Trivedi et al., 2018; Yu et al., 2023). They utilize some form of persistent high-dimensional representation (i.e., a memory) for nodes, which is used to learn the dynamics of the graph as it changes. Some of them also include a GNN-based graph embedding module, which helps in mitigating the problem of memory staleness, i.e., when a node’s memory hasn’t been updated for a long time. We utilize these innovations in our model. Other models that do not involve memory (Xu et al., 2019; Pareja et al., 2020) exist and are covered in a recent survey paper (Longa et al., 2023) on the latest models for learning in dynamic graphs. We highlight that these models do not consider combinatorial optimization problems.
@@
A vast body of machine learning research into solving combinatorial optimization problems exists (Peng et al., 2021; Khalil et al., 2017; Bengio et al., 2021). And, GNNs are the most common backbone for much of this work given their excellent ability to characterize the structural information of graphs (Wu et al., 2021; Zhang et al., 2019). MaxIS is one such problem that has garnered attention from the learning community, and autoregressive approaches (take multiple inference steps to output a solution), such as RL and Dynamic Programming have emerged as a natural choice due to their easy constraint enforcement and a priori integral solution output (Brusca et al., 2024; Ahn et al., 2020). They act as indirect heuristic learning methods. Yet, these methods suffer from sample inefficiency and do not scale well (number of inference steps required scales linearly with problem size) (Peng et al., 2021; Karalias and Loukas, 2020). Non-autoregressive approaches, i.e., single-shot methods (output a solution in one inference step), have been proposed as well (Karalias and Loukas, 2020; Li et al., 2018; Wang et al., 2022; Schuetz et al., 2022) with unsupervised learning being a key attribute of these techniques to mitigate the issue of labeling a vast number of solutions. However, these techniques only investigate static graphs. The only other machine learning model (to the best of our knowledge) that considers combinatorial optimization problems in dynamic graphs is by Gunarathna et al. (2022), and is an RL technique for solving the Traveling Salesman Problem and Vehicle Routing Problem. The most recent and state-of-the-art method for solving MaxIS in static graphs is by Brusca et al. (2024), which is a dynamic programming-based method and suffers from aforementioned drawbacks of autoregressive approaches. Our work is the first single-shot learning method for a combinatorial optimization problem in dynamic graphs.
@@ 3 Maximum Independent Sets in Dynamic Graphs
In general, dynamic graphs (also called dynamic networks, temporal graphs, or evolving graphs) are graphs whose nodes and/or edges change over time (Casteigts et al., 2012; Rossi et al., 2020; Kazemi et al., 2020; Bui-Xuan et al., 2003). For this work, we consider dynamic graphs where at most one edge changes per time and the underlying set of nodes is fixed. Formally, a dynamic graph $\mathcal{G}$ is a finite sequence of graph snapshots $(G_{0},G_{1},\ldots,G_{T})$ where each snapshot $G_{t}=(V,E_{t})$ is an undirected, static graph on nodes $V$ and, for all time steps $0<t\leq T$, we have $|E_{t-1}\oplus E_{t}|=1$, i.e., there is exactly one edge addition or one edge deletion per time step. We denote the edge event at time $t$ which transitions $G_{t-1}$ to $G_{t}$ as $\mathcal{E}_{t}$. The neighborhood of a node $v$ at time $t$ is $\mathcal{N}_{t}(v)=\{u\in V\mid(u,v)\in E_{t}\}$ and its degree is denoted $d_{t}(v)=|\mathcal{N}_{t}(v)|$. The diameter of graph snapshot $G_{t}$, denoted $\text{diam}(G_{t})$, is the maximum hop-distance between any two nodes in $G_{t}$.
We define the Dynamic-Maximum-Independent-Set (Dynamic MaxIS) problem as follows: Given a dynamic graph $\mathcal{G}=(G_{0},G_{1},\ldots,G_{T})$—or, equivalently, an initial graph snapshot $G_{0}$ and a sequence of edge events $(\mathcal{E}_{1},\ldots,\mathcal{E}_{T})$—obtain a MaxIS for each snapshot $G_{t}$. This problem is clearly NP-hard since it contains the NP-complete MaxIS problem as a special case ($T=0$). Given a solution for the MaxIS problem, there exist a polynomial-time conversions to obtain solutions for the Minimum-Vertex-Cover and Maximum-Clique problems; hence, solving Dynamic MaxIS can also aid in solving these problems’ dynamic versions.
@@ 4 An Unsupervised Learning Model for Dynamic MaxIS
As discussed in Section 2, memory-based models have been quite successful in recent years. Motivated by their effectiveness in predictive learning of dynamics, we present an unsupervised learning model focused on the MaxIS problem in dynamic graphs. We take inspiration from the Temporal Graph Networks (TGN) model proposed by Rossi et al. (2020), which generalizes memory-based models by including a GNN-based graph embedding module. We repurpose the two key modules of TGNs, memory and graph embedding, to learn an update mechanism that maintains a candidate MaxIS solution as the underlying graph changes. Moreover, inspired by the use of messages (seen as random perturbations) to improve GNN performance for NP-hard problems (Papp et al., 2021), we couple node memory to an event handling module to inform nodes about edge events.
Our model responds to edge events (i.e., edge additions or deletions) over time. At each time step, the event handling module is responsible for signaling all nodes within an $\alpha$ hop-distance of the edge event that the event occurred. These nodes then use the memory module to update their internal representations using this signal. Finally, each node within a $\beta$ hop-distance of the edge event uses the local aggregation (graph embedding) module to produce a probability of membership in the MaxIS solution (i.e., an estimate) based on its memory, its degree, and the memories of its neighbors. We describe each module and the training-related details of our model in the following subsections.
@@
$s_{t}(v)=[\texttt{enc}(\mathcal{E}_{t})~{}||~{}r_{t}(v)],$
where enc encodes edge deletions as $[0,1]$ and edge additions as $[1,0]$, and $r_{t}(v)$ is the hop-distance from $v$ to $\mathcal{E}_{t}$ normalized by $\alpha$ and linearly interpolated into $[-1,1]$.
@@
$m_{t}(v)=\left\{\begin{array}[]{cl}\texttt{GRU}\big{(}[m_{t-1}(v)~{}||~{}s_{t}(v)],m_{t-1}(v)\big{)}&\text{if $v$ is within an $\alpha$ hop-distance of $\mathcal{E}_{t}$};\\ m_{t-1}(v)&\text{otherwise}.\end{array}\right.$
@@
$\tilde{h}_{t}(v)=\texttt{ReLU}\left(\sum_{u\in\mathcal{N}_{t}(v)}\mathbf{W}_{1}m_{t}(u)\right),$
@@
$\ell_{t}(v)=-p_{t}(v)+\frac{c}{2d_{t}(v)}\sum_{u\in\mathcal{N}_{t}(v)}p_{t}(u)p_{t}(v).$
The $-p_{t}(v)$ term rewards nodes with large estimates (i.e., probabilities of MaxIS membership closer to 1), aligning with the goal of finding a large independent set. This is counterbalanced by the $\sum_{u\in\mathcal{N}_{t}(v)}p_{t}(u)p_{t}(v)$ term which penalizes violations of independence, i.e., when neighbors $u$ and $v$ both have large estimates. This constraint violation term is normalized by the degree of $v$ and is then halved to compensate for double-counting the loss from the perspectives of both $u$ and $v$. Finally, the constant $c$ is a hyperparameter balancing the two terms’ loss contributions; in practice, $c=3$ worked reasonably well across our evaluations (Section 5).
The cumulative loss at time $t$ is the sum of all local losses $\ell_{t}(v)$ for nodes $v$ within a $\beta$ hop-distance of the edge event $\mathcal{E}_{t}$ (i.e., all nodes that updated their estimates at time $t$). This set of nodes at this one time step forms a single batch in our training process. By setting $\beta$ as an appropriate function of dynamic graph size (see Model Variants, below), batch sizes will automatically scale to any input dynamic graph and provide sufficient inference samples for training. The sequential processing of all edge events in a dynamic graph’s training set forms a single epoch in our training process. In a single run, we train until the losses for epochs have stabilized and select the epoch with the least loss.
@@
As explained above, the hop-distances $\alpha$ and $\beta$ control which nodes utilize an edge event to update their memories and which nodes update their estimates and contribute to cumulative loss, respectively. In the bounded cascading (BCAS) model variant, we set $\alpha=\beta=\gamma\cdot\text{diam}(G_{0})$, where $\gamma\ll 1$, strictly bounding nodes’ memory and estimate updates to a local region around each edge event. This variant mimics rule-based approximate update algorithms for MaxIS which analogously spread out from the point of the edge event, updating nodes’ memberships to generate a new MaxIS (Chechik and Zhang, 2021; Behnezhad et al., 2019). Locally bounding these update regions typically yields very fast update operations, at the potential risk of missing better MaxIS solutions that require far-reaching updates.
@@
Pre-Training and Unsupervised Learning.
Many rule-based update algorithms for MaxIS crucially require an existing MaxIS solution for the initial snapshot $G_{0}$ which is later updated as the graph structure changes. We mitigate this (potentially very expensive) prerequisite by utilizing a pre-training phase that efficiently produces memories for all nodes in $G_{0}$ and model weights that training can warm-start from. The idea is to build up $G_{0}$ one edge at a time (as a sequence of edge addition events), updating only the memories of the nodes involved in each edge addition. Once $G_{0}$ is fully constructed, estimates and corresponding losses are computed for all nodes. This constitutes a single epoch of pre-training; in a single pre-training run, we execute epochs until the cumulative loss has stabilized. After executing multiple pre-training runs with random seeds, we extract the node memories and model weights from the run with the least cumulative loss to warm-start model training. Thus, the model as a whole learns to first find a candidate MaxIS (pre-training) and then maintain it (training) in an entirely unsupervised manner.
@@
During testing and deployment, we use a simple rounding and violation removal procedure to convert the model’s relaxed node estimates into a candidate MaxIS. Specifically, all nodes with estimates 0.5 or larger are initially included in the candidate solution; then, nodes violating independence are removed in decreasing order of number of violations until an independent set is obtained. Such a correction/completion procedure is a necessary part of single-shot learning methods that directly use relaxation to solve optimization problems with hard constraints (Schuetz et al., 2022; Donti et al., 2020; Wang et al., 2022; Karalias and Loukas, 2020).

### v2

# Unsupervised Learning of Local Updates for Maximum Independent Set in Dynamic Graphs

@@ Abstract
We present the first unsupervised learning model for finding Maximum Independent Sets (MaxIS) in dynamic graphs where edges change over time. Our method combines structural learning from graph neural networks (GNNs) with a learned distributed update mechanism that, given an edge addition or deletion event, modifies nodes’ internal memories and infers their MaxIS membership in a single, parallel step. We parameterize our model by the update mechanism’s radius and investigate the resulting performance–runtime tradeoffs for various dynamic graph topologies. We evaluate our model against a mixed integer programming solver and the state-of-the-art learning-based methods for MaxIS on static graphs (ICML 2020; NeurIPS 2020, 2023). Across synthetic and empirical dynamic graphs of 50–1,000 nodes, our model achieves competitive approximation ratios with excellent scalability; on large graphs, it significantly outperforms the state-of-the-art learning methods in solution quality, runtime, and memory usage. When generalizing to graphs of 10,000 nodes (100x larger than the ones used for training), our model produces MaxIS solutions 1.05–1.18x larger than any other learning method, even while maintaining competitive runtimes.
@@ 1 Introduction
Combinatorial optimization problems on graphs (e.g., Graph-Coloring, Traveling-Salesman, Vehicle-Routing, etc.) arise naturally in many practical domains (Applegate et al., 2007; Toth and Vigo, 2014). However, many of these problems are classically intractable to solve exactly, and some even resist efficient approximation. The situation is exacerbated by the fact that many real-world applications are dynamic, requiring algorithms to not only find solutions but also maintain them as the underlying structure evolves. While one could iteratively apply algorithms designed for static graphs at each time step, this approach is often computationally expensive and inefficient. Critically, this approach ignores the often close relationship between successive graph structures. A more effective strategy might avoid repeatedly recomputing solutions from scratch by instead leveraging these small structural changes to incrementally update an existing solution. Such update algorithms are an active area of theoretical algorithms research (Assadi et al., 2019; Hanauer et al., 2022; Chechik and Zhang, 2019; Behnezhad et al., 2019; Zheng et al., 2019; Gao et al., 2022), but no learning algorithms have yet taken this approach.
@@
In this paper, we consider the Maximum-Independent-Set (MaxIS) problem in a dynamic setting. For static graphs, MaxIS is classically known to be NP-hard (Garey and Johnson, 1979) and hard to approximate: for general graphs, deterministic approximation within a constant factor is impossible (Robson, 1986) and within an $n^{1-\varepsilon}$ factor is NP-hard (Håstad, 1996). Nevertheless, MaxIS has a diverse range of real-world applications, from identifying functional nodes in brain networks (Afek et al., 2011) to constructing diversified investment portfolios (Hidaka et al., 2023). For example, different stocks can be modeled as nodes and those that are closely related have an edge between them to form a conflict graph; thus, the problem of building a diverse portfolio reduces to finding a high-value independent set. In many of these situations, the conflict graph changes dynamically over time due to factors such as neuronal development, stock market fluctuations, and other unpredictable events. Our goal is to solve MaxIS on such dynamically evolving graphs.
@@
Recently, the learning community has developed several methods for solving combinatorial optimization problems on static graphs (Bengio et al., 2021). Central to many of these methods are Graph Neural Networks (GNNs) due to their efficiency in capturing the structural information present in real-world networks such as the Internet, social networks, and molecular interactions (Zhang et al., 2019; Wu et al., 2020). In the dynamic setting, recent models such as temporal GNNs are primarily concerned with tasks such as edge prediction, node classification, graph classification, etc. (Xu et al., 2020; Rossi et al., 2020). Work considering combinatorial optimization problems in a dynamic setting remains limited, with few examples such as Gunarathna et al. (2022), which is an indirect heuristic learning approach. No existing models learn update mechanisms comparable to theoretical, rule-based update algorithms.
@@
We propose a model that solves MaxIS in dynamic graphs by directly learning an update mechanism instead of a heuristic. We utilize the power of message-passing GNNs (Battaglia et al., 2018) combined with sequence learning to learn an update mechanism that maintains an approximate MaxIS solution as the underlying graph changes. GNNs learn to aggregate structural information and update a node’s membership in the MaxIS solution at any instance of a dynamic graph, while sequence learning learns the effect of a change in the dynamic graph on the change in the MaxIS solution. Having learned an update mechanism, the model is able to update a MaxIS solution following a change in the dynamic graph in a single inference step, unlike more popular heuristic learning methods which require multiple inference steps and thus scale poorly with graph size. We underscore this as a necessary shift in perspective required to address some of the common criticisms of learning approaches to solving combinatorial optimization problems, namely, lack of scalability, computation inefficiency, and huge training sample requirements.
@@
We present the first unsupervised learning model for solving Maximum-Independent-Set (MaxIS) on dynamic graphs. The central idea of learning local, distributed update mechanisms for combinatorial optimization problems on dynamic graphs is novel, and we demonstrate its efficiency in our evaluations. Compared to a commercial mixed integer programming solver (Gurobi Optimization, LLC, 2024), our model maintains favorable approximation ratios across a variety of dynamic graph topologies and sizes. For larger datasets, our model conclusively outperforms the three prior learning methods for MaxIS on static graphs: the current state-of-the-art, DP-GNN (Brusca et al., 2023); a single-shot method, Erdős-GNN (Karalias and Loukas, 2020); and a deep reinforcement learning method, LwDMIS (Ahn et al., 2020). Resource scaling for our method vs. DP-GNN is drastic; we obtain solutions up to 104x faster and use up to 27.5x less memory while producing up to 5.64x larger MaxIS solutions. Compared to Erdős-GNN and LwDMIS, our model delivers 1.071.22x larger MaxIS solutions than either method while also running 6.1824.8x faster and saving up to 2.7812.97x memory. These gains hold even in generalization experiments, where we evaluate our model on graphs 100x the size of those it was trained on.
@@ 2 Related Work
Maximum-Independent-Set (MaxIS) is a classical NP-complete problem in graph theory and has been explored since the 1970s in the field of combinatorial optimization. Algorithms seeking exact solutions use reduction rules (Tarjan and Trojanowski, 1977), branching heuristics (Fomin and Kaski, 2013), and kernelization techniques (Lamm et al., 2016) to reduce the size of the search space. Among exact algorithms, the state-of-the-art is Xiao and Nagamochi (2017) which runs in $1.1996^{n}n^{O(1)}$ time and polynomial space, where $n$ is the number of nodes in the graph. Extensions of MaxIS to dynamic graphs remain relatively unexplored. Zheng et al. (2019) and Gao et al. (2022) have developed update algorithms using rule-based optimization techniques for incrementally maintaining an existing solution in response to edge insertions or deletions. However, no such approaches have been taken by the learning community, which we discuss next.
@@
Memory-based models utilize various forms of persistent, high-dimensional representations (i.e., memories) for nodes to learn the dynamics of a graph as it changes. These have been very successful in efficiently learning prediction and classification tasks in dynamic graphs (Rossi et al., 2020; Kumar et al., 2019; Trivedi et al., 2019; Yu et al., 2023). Some models also include a GNN-based graph embedding module to mitigate the problem of memory staleness, i.e., when a node’s memory hasn’t been updated for a long time. We utilize these innovations in our model. Other learning models for dynamic graphs do not involve memory (Xu et al., 2020; Pareja et al., 2020; Longa et al., 2023) but have never been applied to combinatorial optimization problems.
@@
A vast body of machine learning research considers combinatorial optimization problems (Peng et al., 2021; Dai et al., 2017; Bengio et al., 2021). Among these, GNNs are the most common backbone given their excellent ability to characterize the structural information of graphs (Zhang et al., 2019; Wu et al., 2020). For MaxIS specifically, autoregressive approaches (i.e., those that take multiple inference steps to output a solution) such as reinforcement learning (RL) and dynamic programming (DP) have emerged as natural choices due to their easy constraint enforcement and a priori integral solution output (Brusca et al., 2023; Ahn et al., 2020). Yet, these methods suffer from sample inefficiency poor scalability, requiring a number of inference steps that scales linearly with problem size (Karalias and Loukas, 2020; Peng et al., 2021). Non-autoregressive approaches (i.e., "single-shot" methods that output a solution in one inference step) have also been proposed (Li et al., 2018; Karalias and Loukas, 2020; Wang et al., 2022; Schuetz et al., 2022), typically featuring unsupervised learning to avoid the issue of labeling a vast number of solutions. However, these techniques only investigate static graphs. To the best of our knowledge, the only other learning model that considers combinatorial optimization problems in dynamic graphs is by Gunarathna et al. (2022) applying RL to dynamic versions of the Traveling-Salesman and Vehicle-Routing problems. Our work is the first non-autoregressive, unsupervised learning method for a combinatorial optimization problem in dynamic graphs.
@@ 3 Maximum Independent Sets in Dynamic Graphs
Dynamic graphs (also called dynamic networks, temporal graphs, or evolving graphs) are graphs whose nodes and/or edges change over time (Bui-Xuan et al., 2003; Casteigts et al., 2012; Rossi et al., 2020; Kazemi et al., 2020). For this work, we consider dynamic graphs where at most one edge changes per time and the underlying set of nodes is fixed. Formally, a dynamic graph $\mathcal{G}$ is a finite sequence of graph snapshots $(G_{0},G_{1},\ldots,G_{T})$ where each snapshot $G_{t}=(V,E_{t})$ is an undirected, static graph on nodes $V$ and, for all time steps $0<t\leq T$, we have $|E_{t-1}\oplus E_{t}|=1$, i.e., there is exactly one edge addition or one edge deletion per time step. We denote the edge event at time $t$ which transitions $G_{t-1}$ to $G_{t}$ as $\mathcal{E}_{t}$. The neighborhood of a node $v$ at time $t$ is $\mathcal{N}_{t}(v)=\{u\in V\mid(u,v)\in E_{t}\}$ and its degree is denoted $d_{t}(v)=|\mathcal{N}_{t}(v)|$. The diameter of graph snapshot $G_{t}$, denoted $\text{diam}(G_{t})$, is the maximum hop-distance between any two nodes in $G_{t}$.
We define the Dynamic-Maximum-Independent-Set (Dynamic MaxIS) problem as follows: Given a dynamic graph $\mathcal{G}=(G_{0},G_{1},\ldots,G_{T})$—or, equivalently, an initial graph snapshot $G_{0}$ and a sequence of edge events $(\mathcal{E}_{1},\ldots,\mathcal{E}_{T})$—obtain a MaxIS for each snapshot $G_{t}$. This problem is clearly NP-hard since it contains the NP-complete MaxIS problem as a special case ($T=0$). Given a solution for the MaxIS problem, there exist a polynomial-time conversions to obtain solutions for the Minimum-Vertex-Cover and Maximum-Clique problems; hence, achieving good approximations for Dynamic MaxIS can also aid in approximating these problems’ dynamic versions.
@@ 4 An Unsupervised Learning Model for Dynamic MaxIS
As discussed in Section 2, memory-based models have been quite successful in recent years. Motivated by their effectiveness in predictive learning of dynamics, we present an unsupervised learning model focused on the MaxIS problem in dynamic graphs. We take inspiration from the Temporal Graph Networks (TGN) model proposed by Rossi et al. (2020) which generalizes memory-based models by including a GNN-based graph embedding module. We repurpose the two key modules of TGNs, memory and graph embedding, to learn an update mechanism that maintains a candidate MaxIS solution as the underlying graph changes. Moreover, inspired by the use of messages (seen as random perturbations) to improve GNN performance for NP-hard problems (Papp et al., 2021), we couple node memory to an event handling module to inform nodes about nearby edge events.
Figure 1: An overview of our unsupervised learning model for MaxIS on dynamic graphs. Nodes maintain memories (node colors) which are updated by signals propagating in an $\alpha$-radius from edge events (orange arrows). Each node within a $\beta$-radius of the edge event then locally aggregates its updated memory and those of its neighbors to estimate its new probability of MaxIS membership. These estimates are then used to compute loss (during training) or to generate an integral MaxIS solution (during testing and inference).
Our model responds to edge events (i.e., edge additions or deletions) over time (see Figure 1). At each time step, the event handling module is responsible for signaling all nodes within an $\alpha$ hop-distance of the edge event that the event occurred. These nodes then use the memory module to update their internal representations using this signal. Finally, each node within a $\beta$ hop-distance of the edge event uses the local aggregation (graph embedding) module to produce a probability of membership in the MaxIS solution (i.e., an estimate) based on its memory, its degree, and the memories of its neighbors. We describe each module and the training-related details of our model in the following subsections; pseudocode can be found in Appendix A.
@@
$s_{t}(v)=[\texttt{enc}(\mathcal{E}_{t})~||~r_{t}(v)],$ (1)
where enc encodes edge deletions as $[0,1]$ and edge additions as $[1,0]$, and $r_{t}(v)$ is the hop-distance from $v$ to $\mathcal{E}_{t}$ normalized by $\alpha$ and linearly interpolated into [1,1][-1,1].
@@
[…]
@@
$\tilde{h}_{t}(v)=\texttt{ReLU}\left(\sum_{u\in\mathcal{N}_{t}(v)}\mathbf{W}_{1}m_{t}(u)\right),$ (3)
@@
$\ell_{t}(v)=-p_{t}(v)+\frac{c}{2d_{t}(v)}\sum_{u\in\mathcal{N}_{t}(v)}p_{t}(u)p_{t}(v).$ (6)
The $p_{t}(v)$ term rewards nodes with large estimates (i.e., probabilities of MaxIS membership closer to 1), aligning with the goal of finding a large independent set. This is counterbalanced by the $\sum_{u\in\mathcal{N}_{t}(v)}p_{t}(u)p_{t}(v)$ term which penalizes violations of independence, i.e., when neighbors $u$ and $v$ both have large estimates. This constraint violation term is normalized by the degree of $v$ and is then halved to compensate for double-counting the loss from the perspectives of both $u$ and $v$. Finally, the constant $c$ is a hyperparameter balancing the two terms’ loss contributions; in practice, $c=3$ worked reasonably well across our evaluations (Section 5).
The cumulative loss at time $t$ is the sum of all local losses $\ell_{t}(v)$ for nodes $v$ within a $\beta$ hop-distance of the edge event $\mathcal{E}_{t}$ (i.e., all nodes that updated their estimates at time $t$). This set of nodes at this one time step forms a single batch in our training process. By setting $\beta$ as an appropriate function of dynamic graph size (see Model Variants, below), batch sizes will automatically scale to any input dynamic graph and provide sufficient inference samples for a training step. The sequential processing of all edge events in a dynamic graph’s training set forms a single epoch in our training process. In a single run, we train until the losses for epochs have stabilized and select the epoch with the least loss.
@@
As explained above, the hop-distances $\alpha$ and $\beta$ control which nodes utilize an edge event to update their memories and which nodes update their estimates and contribute to cumulative loss, respectively. In the bounded cascading (BCAS) model variant, we set $\alpha=\beta=\gamma\cdot\text{diam}(G_{0})$, where $\gamma\ll 1$, strictly bounding nodes’ memory and estimate updates to a local region around each edge event. This variant mimics rule-based approximate update algorithms for MaxIS which analogously spread out from the point of the edge event, updating nodes’ memberships to generate a new MaxIS (Behnezhad et al., 2019; Chechik and Zhang, 2019). Locally bounding these update regions typically yields very fast update operations, at the potential risk of missing better MaxIS solutions that require far-reaching updates.
@@
Initial MaxIS Generation and Unsupervised Learning.
Update algorithms for MaxIS crucially require an existing MaxIS solution for the initial snapshot $G_{0}$ which is later updated as the graph structure changes. We mitigate this (potentially very expensive) prerequisite by utilizing an initial MaxIS generation phase that efficiently produces memories for all nodes in $G_{0}$ and model weights that training can warm-start from. The idea is to build up $G_{0}$ one edge at a time (as a sequence of edge addition events), updating only the memories of the nodes involved in each edge addition. Once $G_{0}$ is fully constructed, estimates and corresponding losses are computed for all nodes. This constitutes a single epoch of this generation phase; in a single run of the generation phase, we execute epochs until the cumulative loss has stabilized. After executing multiple runs with random seeds, we extract the node memories and model weights from the run with the least cumulative loss to warm-start model training. Thus, the model as a whole learns to first find a candidate MaxIS (generation phase) and then maintain it (training phase) in an entirely unsupervised manner.
@@
During testing and deployment, we use a simple rounding and violation removal procedure to convert the model’s relaxed node estimates into a candidate MaxIS. Specifically, all nodes with estimates 0.5 or larger are initially included in the candidate solution; then, nodes violating independence are removed in decreasing order of number of violations until an independent set is obtained. Such a procedure is a necessary part of single-shot learning methods that directly use relaxation to solve optimization problems with hard constraints (Karalias and Loukas, 2020; Donti et al., 2021; Schuetz et al., 2022; Wang et al., 2022).